# A Two-Timescale Primal-Dual Framework for Reinforcement Learning via Online Dual Variable Guidance

### v2

@@ C.3 Auxiliary Results and Proofs
$T_{5}$ $=\mathbb{E}[\langle z_{k-\tau_{k}},\mathbb{E}[\tilde{g}_{k+1-\tau_{k}}(\rho_{k-\tau_{k}},V_{k-\tau_{k}};X_{k})|\mathcal{F}_{k-\tau_{k}}^{-}]-\mathbb{E}[\tilde{g}_{k+1-\tau_{k}}(\rho_{k-\tau_{k}},V_{k-\tau_{k}};\widetilde{X}_{k})|\mathcal{F}_{k-\tau_{k}}^{-}]\rangle|\mathcal{G}_{\delta}]$
$T_{6}$ $=\mathbb{E}[\langle z_{k-\tau_{k}},\mathbb{E}[\tilde{g}_{k+1-\tau_{k}}(\rho_{k-\tau_{k}},V_{k-\tau_{k}};\widetilde{X}_{k})|\mathcal{F}_{k-\tau_{k}}^{-}]-\mathbb{E}_{\bar{X}\sim\mu_{\rho_{k-\tau_{k}}}}[\tilde{g}_{k+1-\tau_{k}}(\rho_{k-\tau_{k}},V_{k-\tau_{k}};\bar{X})]\rangle|\mathcal{G}_{\delta}]$
$T_{7}$ $=\mathbb{E}[\langle z_{k-\tau_{k}},\mathbb{E}_{\bar{X}\sim\mu_{\rho_{k-\tau_{k}}}}[\tilde{g}_{k+1-\tau_{k}}(\rho_{k-\tau_{k}},V_{k-\tau_{k}};\bar{X})]-\mathbb{E}_{X\sim\mu_{\rho_{k}}}[\tilde{g}_{k+1-\tau_{k}}(\rho_{k-\tau_{k}},V_{k-\tau_{k}};X)]\rangle|\mathcal{G}_{\delta}]$
@@
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +\mathbb{E}[\langle\lambda(\rho_{k-\tau_{k}})-\lambda(\rho_{k}),\mathbb{E}\big[\widetilde{G}_{k+1}|\mathcal{F}_{k}^{-}\big]-\mathbb{E}_{X\sim\mu_{\rho_{k}}}[\tilde{g}_{k+1}(\rho_{k},V_{k};X)]\rangle|\mathcal{G}_{\delta}]$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +\mathbb{E}[\langle z_{k-\tau_{k}},\mathbb{E}\big[\widetilde{G}_{k+1}|\mathcal{F}_{k}^{-}\big]-\mathbb{E}_{X\sim\mu_{\rho_{k}}}[\tilde{g}_{k+1}(\rho_{k},V_{k};X)]\rangle|\mathcal{G}_{\delta}]$
@@
The term $T_{4}$ addresses the drift in the buffer over the mixing interval. Using the linearity of the gradient $\nabla_{V}L$ in the transition probabilities (see (32a)), we have
@@
where for the last inequality we recall the bound $\|\rho\|_{1}\leq|\mathcal{X}|C^{U}$ for all $\rho\in H$ from Lemma C.2. Fix a state-action pair $x=(s,a)\in\mathcal{X}$, and set $m:=\nu_{k-\tau_{k}}(x)$. Let $v_{x}\leq\tau_{k}$ be the number of new samples for $x$ in the window, and let $u_{s}$ be the number of transitions from $x$ to $s$ in the window. Let $c_{s}\in\mathbb{N}$ be such that $\mathcal{P}_{\text{old}}(s|x):=\frac{c_{s}}{m}$ and $\mathcal{P}_{\text{new}}(s|x):=\frac{c_{s}+u_{s}}{m+v_{x}}$ are the empirical next-state distributions, then
@@
To bound the stationary law drift comparison term $T_{7}$, we use the Lipschitz continuity of the stationary law, derived in Lemma C.7, and obtain
@@
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ -\mathbb{E}_{X\sim\mu_{\rho_{k}}}[\tilde{g}_{k+1-\tau_{k}}(\rho_{k-\tau_{k}},V_{k-\tau_{k}};X)]\|_{2}|\mathcal{G}_{\delta}]$
@@
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +2D(LB+D)\tau_{k}\alpha_{\mathrm{env},k-\tau_{k}}$
@@ Appendix D Additional Proofs
Proof of Corollary 2.1
The proof for the entropy-regularized MDP is provided in [Ying and Zhu, 2020, Section 3]. We adapt the proof to apply to a general strongly convex regularizer. Although the proof closely follows Ying and Zhu [2020], we include it to ensure the paper remains self-contained. Throughout this proof we adopt the shorthand notation $v_{s}$ or $v_{sa}$ for $v(s)$ and $v(s,a)$, respectively for vectors in $\mathbb{R}^{|\mathcal{S}|}$ and $\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$ or corresponding mappings from $\mathcal{S}$ or $\mathcal{S}\times\mathcal{A}$ to $\mathbb{R}$.
@@
Inserting the decomposition above into (7) and using the convex-concave structure of the regularized Lagrangian, which allows us to use Sion’s minmax theorem [Sion, 1958], we obtain
@@
The constraint of the primal LP is equivalently expressed as $T_{\star,G}V\leq V$ with the regularized Bellman optimality operator defined in (6). Let $\widehat{V}$ be a solution to $P_{r}$. Then by primal feasibility, the slack variable $S:=\widehat{V}-T_{\star,G}\widehat{V}\geq 0$. Assume $S\neq 0$. Note that $T_{\star,G}$ is a monotone operator, [Geist et al., 2019, Prop. 2], that is, for $V_{1}\geq V_{2}$, $T_{\star,G}V_{1}\geq T_{\star,G}V_{2}$. Therefore $\widetilde{V}:=T_{\star,G}\widehat{V}$ fulfills $T_{\star,G}\widetilde{V}\leq T_{\star,G}\widehat{V}=\widetilde{V}$, that is, $\widetilde{V}$ is primal feasible. For $S\neq 0$, there exists a state $s$ with $\widetilde{V}_{s}<\widehat{V}_{s}$ and hence $\mu^{\top}\widetilde{V}<\mu^{\top}\widehat{V}$. This contradicts the optimality of $\widehat{V}$, therefore $S=0$ holds. The solution to the primal problem is the unique fixed point of the regularized Bellman optimality operator.
@@
Proof of Lemma 2.1
We begin by establishing an upper bound for $\rho^{\star}$. The optimal primal and dual solutions are linked by the first-order optimality condition of the primal variable $V$. We then use the bound on $V^{\star}$ to obtain a bound on $\rho^{\star}$. Moreover, thanks to Theorem 2.3, we know that the tuple $(\pi^{\star}$, $V^{\star})$ is a fixed point of the regularized Bellman operator (6) and therefore fulfills the regularized Bellman equation
@@
Define $\Delta_{\max}(s):=\max_{a\in\mathcal{A}}\Delta[V^{\star}]_{sa}$ and $\Delta_{\min}(s):=\min_{a\in\mathcal{A}}\Delta[V^{\star}]_{sa}$. We then denote the maximum difference as $\Delta_{s}:=\Delta_{\max}(s)-\Delta_{\min}(s)$. Set $K_{1}:=|\mathcal{A}|-1$, $K_{2}(\Delta_{s}):=\exp(-\frac{1}{\eta_{\rho}}\Delta_{s})$. The minimum entropy of the softmax policy is obtained when in state $s$ one action gets the “high" weight, $\Delta_{\max}(s)$, and the other $K_{1}$ states get “low" weight $\Delta_{\min}$. In that case, the policy assigns the probabilities

### v3

@@ C.3 Auxiliary Results and Proofs
$T_{5}$ $=$
$\mathbb{E}[\langle z_{k-\tau_{k}},\mathbb{E}[\tilde{g}_{k+1-\tau_{k}}(\rho_{k-\tau_{k}},V_{k-\tau_{k}};X_{k})|\mathcal{F}_{k-\tau_{k}}^{-}]-\mathbb{E}[\tilde{g}_{k+1-\tau_{k}}(\rho_{k-\tau_{k}},V_{k-\tau_{k}};\widetilde{X}_{k})|\mathcal{F}_{k-\tau_{k}}^{-}]\rangle|\mathcal{G}_{\delta}]$
$T_{6}$ $=$
$\mathbb{E}[\langle z_{k-\tau_{k}},\mathbb{E}[\tilde{g}_{k+1-\tau_{k}}(\rho_{k-\tau_{k}},V_{k-\tau_{k}};\widetilde{X}_{k})|\mathcal{F}_{k-\tau_{k}}^{-}]-\mathbb{E}_{\bar{X}\sim\mu_{\rho_{k-\tau_{k}}}}[\tilde{g}_{k+1-\tau_{k}}(\rho_{k-\tau_{k}},V_{k-\tau_{k}};\bar{X})]\rangle|\mathcal{G}_{\delta}]$
$T_{7}$ $=$
$\mathbb{E}[\langle z_{k-\tau_{k}},\mathbb{E}_{\bar{X}\sim\mu_{\rho_{k-\tau_{k}}}}[\tilde{g}_{k+1-\tau_{k}}(\rho_{k-\tau_{k}},V_{k-\tau_{k}};\bar{X})]-\mathbb{E}_{X\sim\mu_{\rho_{k}}}[\tilde{g}_{k+1-\tau_{k}}(\rho_{k-\tau_{k}},V_{k-\tau_{k}};X)]\rangle|\mathcal{G}_{\delta}]$
@@
$~~~~~+\mathbb{E}[\langle\lambda(\rho_{k-\tau_{k}})-\lambda(\rho_{k}),\mathbb{E}\big[\widetilde{G}_{k+1}|\mathcal{F}_{k}^{-}\big]-\mathbb{E}_{X\sim\mu_{\rho_{k}}}[\tilde{g}_{k+1}(\rho_{k},V_{k};X)]\rangle|\mathcal{G}_{\delta}]$
$~~~~~+\mathbb{E}[\langle z_{k-\tau_{k}},\mathbb{E}\big[\widetilde{G}_{k+1}|\mathcal{F}_{k}^{-}\big]-\mathbb{E}_{X\sim\mu_{\rho_{k}}}[\tilde{g}_{k+1}(\rho_{k},V_{k};X)]\rangle|\mathcal{G}_{\delta}]$
@@
The term $T_{4}$ addresses the drift in the buffer over the mixing interval. Using the linearity of the gradient $\nabla_{V}L$ in the transition probabilities (see 32a), we have
@@
where for the last inequality we recall the bound $\|\rho\|_{1}\leq|\mathcal{X}|C^{U}$ for all $\rho\in H$ from Lemma 26. Fix a state-action pair $x=(s,a)\in\mathcal{X}$, and set $m:=\nu_{k-\tau_{k}}(x)$. Let $v_{x}\leq\tau_{k}$ be the number of new samples for $x$ in the window, and let $u_{s}$ be the number of transitions from $x$ to $s$ in the window. Let $c_{s}\in\mathbb{N}$ be such that $\mathcal{P}_{\text{old}}(s|x):=\frac{c_{s}}{m}$ and $\mathcal{P}_{\text{new}}(s|x):=\frac{c_{s}+u_{s}}{m+v_{x}}$ are the empirical next-state distributions, then
@@
To bound the stationary law drift comparison term $T_{7}$, we use the Lipschitz continuity of the stationary law, derived in Lemma 32, and obtain
@@
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~-\mathbb{E}_{X\sim\mu_{\rho_{k}}}[\tilde{g}_{k+1-\tau_{k}}(\rho_{k-\tau_{k}},V_{k-\tau_{k}};X)]\|_{2}|\mathcal{G}_{\delta}]$
@@
$~~~~~~~~~+2D(LB+D)\tau_{k}\alpha_{\mathrm{env},k-\tau_{k}}$
@@
The bound on the Markov mismatch term in Proposition 19 is given by (Zeng et al., 2024, Lemma 1) which is restated here for completeness, without proof. We adapted the statement to our notation using $\bar{\beta}_{k}\leq\beta_{\mathrm{env},k}$ on $\mathcal{G}_{\delta}$.
Lemma 34
Fix $\delta\in(0,1)$ and let $k\geq\max\{K(\delta),\mathcal{K}\}$. On the event $\mathcal{G}_{\delta}$
$\mathbb{E}\!\left[\left\langle\nabla f(\rho_{k}),\,\Delta^{X}_{k+1}\right\rangle\;\middle|\;\mathcal{G}_{\delta}\right]\leq 12L^{2}B^{3}\,\tau_{k}^{2}\,\beta_{\mathrm{env},k-\tau_{k}}.$
@@ Appendix D Additional Proofs
Proof of Corollary 2
The proof for the entropy-regularized MDP is provided in (Ying and Zhu, 2020, Section 3). We adapt the proof to apply to a general strongly convex regularizer. Although the proof closely follows Ying and Zhu (2020), we include it to ensure the paper remains self-contained. Throughout this proof we adopt the shorthand notation $v_{s}$ or $v_{sa}$ for $v(s)$ and $v(s,a)$, respectively for vectors in $\mathbb{R}^{|\mathcal{S}|}$ and $\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$ or corresponding mappings from $\mathcal{S}$ or $\mathcal{S}\times\mathcal{A}$ to $\mathbb{R}$.
@@
Inserting the decomposition above into (7) and using the convex-concave structure of the regularized Lagrangian, which allows us to use Sion’s minimax theorem (Sion, 1958), we obtain
@@
The constraint of the primal LP is equivalently expressed as $T_{\star,G}V\leq V$ with the regularized Bellman optimality operator defined in (6). Let $\widehat{V}$ be a solution to $P_{r}$. Then by primal feasibility, the slack variable $S:=\widehat{V}-T_{\star,G}\widehat{V}\geq 0$. Assume $S\neq 0$. Note that $T_{\star,G}$ is a monotone operator, (Geist et al., 2019, Prop. 2), that is, for $V_{1}\geq V_{2}$, $T_{\star,G}V_{1}\geq T_{\star,G}V_{2}$. Therefore $\widetilde{V}:=T_{\star,G}\widehat{V}$ fulfills $T_{\star,G}\widetilde{V}\leq T_{\star,G}\widehat{V}=\widetilde{V}$, that is, $\widetilde{V}$ is primal feasible. For $S\neq 0$, there exists a state $s$ with $\widetilde{V}_{s}<\widehat{V}_{s}$ and hence $\mu^{\top}\widetilde{V}<\mu^{\top}\widehat{V}$. This contradicts the optimality of $\widehat{V}$, therefore $S=0$ holds. The solution to the primal problem is the unique fixed point of the regularized Bellman optimality operator.
@@
Proof of Lemma 5
We begin by establishing an upper bound for $\rho^{\star}$. The optimal primal and dual solutions are linked by the first-order optimality condition of the primal variable $V$. We then use the bound on $V^{\star}$ to obtain a bound on $\rho^{\star}$. Moreover, thanks to Theorem 4, we know that the tuple $(\pi^{\star}$, $V^{\star})$ is a fixed point of the regularized Bellman operator (6) and therefore fulfills the regularized Bellman equation
@@
Define $\Delta_{\max}(s):=\max_{a\in\mathcal{A}}\Delta[V^{\star}]_{sa}$ and $\Delta_{\min}(s):=\min_{a\in\mathcal{A}}\Delta[V^{\star}]_{sa}$. We then denote the maximum difference as $\Delta_{s}:=\Delta_{\max}(s)-\Delta_{\min}(s)$. Set $K_{1}:=|\mathcal{A}|-1$, $K_{2}(\Delta_{s}):=\exp(-\frac{1}{\eta_{\rho}}\Delta_{s})$. The minimum entropy of the softmax policy is obtained when in state $s$ one action gets the “high” weight, $\Delta_{\max}(s)$, and the other $K_{1}$ states get “low” weight $\Delta_{\min}$. In that case, the policy assigns the probabilities